# Superquantile-Gibbs Relaxation for Minima-selection in Bilevel Optimization

### v2

@@ A.1 Riemannian notions
In this subsection we fix notation for basic Riemannian objects and record definitions that will be used repeatedly in later proofs. Any compact manifold $\mathcal{S}$ of class $\mathcal{C}^{3}$ (indeed $\mathcal{C}^{1}$) without boundary admits a Riemannian metric. [^22] A fundamental theorem in differential geometry states that every paracompact manifold $M$ of class $\mathcal{C}^{k}$ with $k\geq 1$ can be endowed with a Riemannian metric. Given a Riemannian metric $\mathsf{g}$ on $\mathcal{S}$, define
@@
Section˜2 requires that, for each $\theta\in\Theta$, the second fundamental form $\mathrm{I\!I}_{S(\theta)}$ is uniformly bounded in this sense, with a constant $\mathsf{C}>0$ independent of $\theta$. For further background, see also [gong2024poincare, Appendix D].
@@
A.2 Discussion on the validity of assumptions in [hasenpflug2024wasserstein, Theorem 3.6]
The assumptions required by hasenpflug2024wasserstein follow from our standing conditions. First, their positive definiteness requirement for $g(\theta,\cdot)$ holds because the condition implies that $g(\theta,\cdot)$ is locally Morse–Bott [rebjock2023fast, Section 2.3], and hence the Hessian of $g$ is positive definite in directions normal to the minimizer manifold $\mathcal{S}(\theta)$. Second, their tail condition [hasenpflug2024wasserstein, Assumption 3.1] is ensured by our coercivity assumption on $g$.
Finally, Sections˜2 and 2 verify the geometric conditions in [hasenpflug2024wasserstein, Assumption 2.2] (smooth, boundaryless, and connected minimizer manifolds). It is worth noting that the orientability of the manifold is not necessary in our setting. We define the volume measure $\mathcal{M}$ in a coordinate-free manner based on the Riemannian metric $g_{\mathcal{S}}$. Orientability, a stronger condition, ensures the existence of a global top-form rather than just a density, which is not required in our framework.
@@ Appendix B Proofs of results in Section˜3
Consider the decreasing sublevel sets $E_{\varepsilon}:=\{x:\ g(\theta,x)-g_{\star}\leq\varepsilon\}$; if none is contained in $\mathcal{U}(\theta)$, compactness yields a limit point in $\mathcal{S}(\theta)\cap(\mathbb{R}^{d}\setminus\mathcal{U}(\theta))$, a contradiction.
@@
Fix $\theta\in\Theta$. By [rebjock2023fast, Corollary 2.13] and the path-connectedness of $\mathcal{S}(\theta)$, the rank of $\partial_{x}^{2}g(\theta,x)$ is constant over $x\in\mathcal{S}(\theta)$. Denote it by
@@
so that $\texttt{k}(\theta)=d-\mathrm{r}(\theta)$.
@@
Moreover, implies a local Morse–Bott structure for $g(\theta,\cdot)$ [rebjock2023fast, Corollary 2.13], hence the nonzero eigenvalues of $\partial_{x}^{2}g(\bar{\theta},x)$ are bounded away from 0 uniformly for $x$ near $\mathcal{S}(\bar{\theta})$. Therefore, for $n$ large enough, the perturbation above cannot change the number of nonzero eigenvalues, and thus
@@
Lemma B.3 ([bishop2011geometry], Section 9.3).
@@
Proposition B.4 ([petersen2006riemannian], Lemma 6.4.7).
@@
For each fixed $\theta$, $\mathcal{S}(\theta)$ is a compact k-dimensional Riemannian manifold, hence $0<\mathcal{M}(\mathcal{S}(\theta))<\infty$ by the standard coordinate formula for the Riemannian volume element; see [lee2006riemannian, Section “Riemannian volume”].
@@
To establish the boundedness of the density $\rho_{\theta}$, it suffices to show that $|\partial_{r}^{2}g(\theta,y(u,r))|$ is bounded. Since the condition ensures that $g(\theta,\cdot)$ is locally Morse-Bott (see [rebjock2023fast, section 2.3]), we have $\partial_{r}^{2}g(\theta,y(u,r))\big|_{r=0}\succeq\mu\cdot I_{(d-\texttt{k})\times(d-\texttt{k})}$.
@@ Appendix C Proofs of results in Section˜4
$d\mathcal{M}(x)\;=\;\sqrt{\det\bigl(\mathsf{g}_{ij}(x)\bigr)}\,dx^{1}\cdots dx^{\texttt{k}},$
@@
$\int_{U}f(x)\,d\mathcal{M}(x)\;=\;\int_{U}f(x)\,\sqrt{\det\bigl(\mathsf{g}_{ij}(x)\bigr)}\,dx^{1}\cdots dx^{\texttt{k}},$
@@
Fix a point $p\in\mathcal{S}$ and $0<r<r_{0}$. By [petersen2006riemannian, Theorem 5.4.2], the exponential map
@@
$\mathcal{M}\bigl(\mathbb{B}_{\mathsf{g}}(p;r)\bigr)\;=\;\int_{\mathbb{B}_{\mathsf{g}}(p;r)}d\mathcal{M}(x)\;=\;\int_{\mathbb{B}_{\texttt{k}}(0;r)}\bigl(\exp_{p}^{*}d\mathcal{M}\bigr)(v),$ (39)
where $v=(v^{1},\dots,v^{\texttt{k}})$ are coordinates on $T_{p}\mathcal{S}\cong\mathbb{R}^{\texttt{k}}$ and $dv$ is the standard Lebesgue measure on $\mathbb{R}^{\texttt{k}}$.
In normal coordinates at $p$, the exponential map $\exp_{p}$ is precisely the coordinate map, so $\exp_{p}^{*}d\mathcal{M}$ has the expression
$(\exp_{p}^{*}d\mathcal{M})(v)\;=\;\sqrt{\det\bigl(\mathsf{g}_{ij}(v)\bigr)}\,dv^{1}\cdots dv^{\texttt{k}},$
@@
$\mathcal{M}\bigl(\mathbb{B}_{\mathsf{g}}(p;r)\bigr)\;=\;\int_{\mathbb{B}_{\texttt{k}}(0;r)}\sqrt{\det\!\bigl(\mathsf{g}_{ij}(v)\bigr)}\;dv.$ (40)
Step 2: Quadratic control of the metric in normal coordinates, uniformly in $p$. In a normal coordinate neighborhood of $p$, the metric admits the expansion (see, e.g., [petersen2006riemannian, Section 5.4])
@@
$\frac{3}{4}\,I\preceq\mathsf{g}(v)\preceq\frac{5}{4}\,I\quad\Longrightarrow\quad\Bigl(\frac{3}{4}\Bigr)^{\texttt{k}/2}\leq\sqrt{\det\bigl(\mathsf{g}_{ij}(v)\bigr)}\leq\Bigl(\frac{5}{4}\Bigr)^{\texttt{k}/2},\qquad\text{for all }\|v\|\leq r_{*}.$ (44)
@@
$\Bigl(\frac{3}{4}\Bigr)^{\texttt{k}/2}{\texttt{Leb}}_{\texttt{k}}\bigl(\mathbb{B}_{\texttt{k}}(p;r)\bigr)\leq\mathcal{M}\bigl(\mathbb{B}_{\mathsf{g}}(p;r)\bigr)\;=\;\int_{\mathbb{B}_{\texttt{k}}(0;r)}\sqrt{\det\bigl(\mathsf{g}_{ij}(v)\bigr)}\,dv\leq\Bigl(\frac{5}{4}\Bigr)^{\texttt{k}/2}{\texttt{Leb}}_{\texttt{k}}\bigl(\mathbb{B}_{\texttt{k}}(p;r)\bigr).$
@@
where (a) follow by (45). Then, taking the minimum over $\beta$ at the r.h.s. of (46), we get $\min_{\beta^{\prime}}\phi_{0,\delta}(\theta,\beta^{\prime})-\min_{\beta}\phi_{\lambda,\delta}(\theta,\beta)\leq\frac{L_{f,1}}{\delta}(\mu^{\lambda}_{g}(\theta),\mu^{0}_{g}(\theta)))$. Interchanging $\min_{\beta}\phi_{\lambda,\delta}(\theta,\beta)$ and $\min_{\beta^{\prime}}\phi_{0,\delta}(\theta,\beta^{\prime})$, one obtains
@@ Appendix D Proofs of results in Section˜5
$\int_{x:\|x\|>R}\|x\|^{n}e^{-c\|x\|^{2}}dx$
$=\int_{0}^{\pi}\cdots\int_{0}^{\pi}\int_{0}^{2\pi}\int_{r>R}e^{-cr^{2}}r^{d+n-1}\sin^{d-2}(\phi_{1})\sin^{d-3}(\phi_{2})\ldots\sin(\phi_{d-2})drd\phi_{1}\cdots d\phi_{d-1}$
$\leq\int_{0}^{\pi}\cdots\int_{0}^{\pi}\int_{0}^{2\pi}\int_{r>R}e^{-cr^{2}}r^{d+n-1}drd\phi_{1}\cdots d\phi_{d-1}\leq 2\pi^{d-1}\int_{R}^{+\infty}e^{-cr^{2}}r^{d+n-1}dr$
$\leq 2\pi^{d+n-1}c^{-\frac{d+n}{2}}\cdot\Gamma\left(\frac{d+n}{2},cR^{2}\right),$ (48)
where in the last inequality we use the definition of the upper incomplete Gamma function, i.e., $\Gamma(s,x):=\int_{x}^{+\infty}t^{s-1}e^{-t}dt$. By Gautschi’s inequality [^25] By Gautschi-type’s inequality, $\Gamma(x,s)\leq sx^{s-1}e^{-x}$ for $x>s$ and $s\geq 1$. Note that for $s=1$, $\Gamma(s,x)=e^{-x}$., for $cR^{2}>(d+n)/2$ and $d+n\geq 2$,
$\int_{x:\|x\|>R}\|x\|^{n}e^{-c\|x\|^{2}}dx\leq 2\pi^{d+n-1}\cdot c^{-\frac{d+n}{2}}\max\{(d+n)/2,1\}(cR^{2})^{\frac{d+n}{2}-1}e^{-cR^{2}}$
@@
$\mathbb{E}_{X\sim\mu_{\theta}^{\lambda}}[\|X\|^{n}\cdot\bm{1}\left[X\notin\mathbb{B}_{d}(0;2\mathsf{D})\right]]=\frac{1}{Z_{\lambda}}\int_{\|x\|\geq 2\mathsf{D}}\|x\|^{n}e^{-g(x)/\lambda}dx,$
@@
$\mathbb{E}_{X\sim\mu_{\theta}^{\lambda}}[\|X\|^{n}\cdot\bm{1}\left[X\notin\mathbb{B}_{d}(0;2\mathsf{D})\right]]\leq\frac{1}{Z_{\lambda}}\int_{\|x\|\geq 2\mathsf{D}}\|x\|^{n}e^{-\frac{\mu_{\mathsf{qg}}(\|x\|-\mathsf{D})^{2}}{2\lambda}}dx.$
@@
$\mathbb{E}_{X\sim\mu_{\theta}^{\lambda}}[\|X\|^{n}\cdot\bm{1}\left[X\notin\mathbb{B}_{d}(0;2\mathsf{D})\right]]\leq\frac{1}{Z_{\lambda}}\cdot 2\pi^{d+n-1}\cdot\left(\frac{2\lambda}{\mu_{\mathsf{qg}}}\right)\cdot\max\{(d+n)/2,1\}\cdot\mathsf{D}^{d+n-2}e^{-\frac{\mathsf{D}^{2}}{\lambda}}.$ (51)
For a lower bound on $Z_{\lambda}$, we use the estimate (see [hasenpflug2024wasserstein, Remark A.21]): for a compact set $U\subset\mathbb{R}^{d}$ and $x_{0}\in\mathcal{S}$,
$\int_{U}e^{-g(x)/\lambda}dx\geq\lambda^{\frac{d}{2}}\left[(2\pi)^{d/2}\det\left(\frac{\partial^{2}g(x_{0})}{\partial x^{2}}\right)^{-\frac{1}{2}}-c\lambda\right].$
@@
$\frac{1}{2}<\mathbb{P}_{X\sim\mu_{\theta}^{\lambda}}[f(\theta,X)<\beta^{*}(\theta)]$
$\leq\mathbb{P}[f(\theta,X)<\beta^{*}(\theta),X\in\mathbb{B}_{d}(0;2\mathsf{D})]+\mathbb{P}[f(\theta,X)<\beta^{*}(\theta),X\notin\mathbb{B}_{d}(0;2\mathsf{D})]$
@@
Given that (i) $\phi_{\lambda,\delta}(\theta,\cdot)$ is convex and Lipschitz continuous, (ii) the feasible set is bounded, and (iii) the stochastic subgradient $\tilde{\partial}_{\beta}\phi_{\lambda,\delta}(\theta,\beta;X)$ satisfies $|\tilde{\partial}_{\beta}\phi_{\lambda,\delta}(\theta,\beta;X)|\leq\sigma:=1+\delta^{-1}$, we invoke the convergence guarantees for convex Lipschitz objectives [nemirovskij1983problem, lan2012validation] together with the large-deviation analysis for stochastic approximation [juditsky2008large, Thm. 2.1]. Accordingly, the output $\hat{\beta}$ of PSGD (see Algorithm˜1) satisfies
@@
after at most $\mathcal{O}(\mathrm{B}^{2}\sigma^{2}\epsilon_{v}^{-2}\log(\Delta^{-1}))$ iterations.
@@
$L\;\geq\;C\cdot\frac{\mathrm{B}^{2}}{\delta^{2}\epsilon_{v}^{2}}\cdot\log\left(\frac{1}{\epsilon_{v}\delta}\right).$
@@
By picking $\delta=\epsilon_{v}^{\texttt{k}}$ and $|I|\geq 4\tilde{C}_{f}\epsilon_{v}^{-2\texttt{k}}$ and
$L\;\geq\;\mathcal{O}\left(\texttt{k}\cdot\frac{\mathrm{B}^{2}}{\epsilon_{v}^{2(\texttt{k}+1)}}\cdot\log\left(\frac{1}{\epsilon_{v}}\right)\right),$
@@
Step 1: Standard zeroth-order analysis for the smoothed problem. We define the smoothed objective $\psi_{\rho}(\theta):=\mathbb{E}_{u\sim\text{uniform}(\mathbb{S}^{m-1})}[F_{\max}(\theta+\rho u)]$, where $u$ is uniformly distributed on the unit sphere $\mathbb{S}^{m-1}$. As shown in [nesterov2017random, Theorem 1 and Lemma 2], for any $L_{\mathsf{f},1}$-Lipschitz function $\mathsf{f}:\mathbb{R}^{m}\to\mathbb{R}$, the smoothed version $\mathsf{f}_{\rho}(\theta):=\mathbb{E}_{u\sim\text{uniform}(\mathbb{S}^{m-1})}[\mathsf{f}(\theta+\rho u)]$ satisfies:
@@
    $\nabla\mathsf{f}_{\rho}(\theta)\in\partial_{\rho}\mathsf{f}(\theta)$, the $\rho$-subdifferential of $\mathsf{f}$, as shown in [lin2022gradient, Theorem 3.1].
In our scenario, by using Section˜3.1, $F_{\max}$ is Lipschitz continuous with the constant $L_{F_{\max},1}:=L_{f,1}(1+L_{g,2}/\mu)$. By applying the aforementioned findings on $F_{\max}$, it follows that $\psi_{\rho}$ is $L_{\psi,2}=\sqrt{m}L_{F_{\max},1}\rho^{-1}$-smooth. Furthermore, the gradient $\nabla\psi_{\rho}(\theta)$ is contained in the $\rho$-subdifferential set of $F_{\max}$ at the point $\theta$. Therefore, if a point $\theta$ satisfies $\|\mathcal{G}_{\Theta}(\theta,\nabla\psi_{\rho};\eta)\|\leq\epsilon$, then $\theta$ is an $(\epsilon,\rho,\eta)$-generalized Goldstein stationary point of $F_{\max}$ (see Section˜5.2). In what follows, we analyze the convergence of Algorithm 2 to a critical point of $\psi_{\rho}$ in terms of the gradient mapping.

### v4

@@ A.1 Riemannian Notions
In this subsection we fix notation for basic Riemannian objects and record definitions that will be used repeatedly in later proofs. Any compact manifold $\mathcal{S}$ of class $\mathcal{C}^{3}$ (indeed $\mathcal{C}^{1}$) without boundary admits a Riemannian metric [^23] A fundamental theorem in differential geometry states that every paracompact manifold $M$ of class $\mathcal{C}^{k}$ with $k\geq 1$ can be endowed with a Riemannian metric.. Given a Riemannian metric $\mathsf{g}$ on $\mathcal{S}$, define
@@
Section˜2 requires that, for each $\theta\in\Theta$, the second fundamental form $\mathrm{I\!I}_{S(\theta)}$ is uniformly bounded in this sense, with a constant $\mathsf{C}>0$ independent of $\theta$. For further background, see also [35, Appendix D].
@@
A.2 Differentiability of Hyper-objective
The authors of [50, 14] reformulate BLO via the classical value-function approach and assume that a PŁ-type condition (or an error bound) holds for the penalized objective $\sigma f+g$, uniformly over $\sigma\in[0,\sigma_{0}]$ where $\sigma_{0}>0$ is a constant. More precisely, Kwon et al. [50] impose a proximal error bound for $\sigma f+g$, whereas Chen et al. [14] assume a global PŁ condition for $\sigma f+g$; Chen et al. [14, Proposition C.1] show these two assumptions are equivalent.
Under this global regularity, both works establish a Danskin-type result implying that the hyper-objective $F_{\max}$ is differentiable. However, their conclusion also entails that every choice $x\in\mathcal{S}(\theta)$ yields the same hyper-gradient $\nabla F_{\max}(\theta)$, so minima-selection becomes redundant. Finally, requiring a uniform PŁ condition (or error bound) for $\sigma f+g$ is quite stringent and may fail even when $f$ and $g$ individually satisfy PŁ-type properties.
A.3 Discussion on The Validity of Assumptions in [38, Theorem 3.6]
The assumptions required by [38, Theorem 3.6] follow from our standing conditions. First, their positive definiteness requirement for $g(\theta,\cdot)$ holds because the condition implies that $g(\theta,\cdot)$ is locally Morse–Bott [73, Section 2.3], and hence the Hessian of $g$ is positive definite in directions normal to the minimizer manifold $\mathcal{S}(\theta)$. Second, their tail condition [38, Assumption 3.1] is ensured by our coercivity assumption on $g$.
Finally, Sections˜2 and 2 verify the geometric conditions in [38, Assumption 2.2] (smooth, boundaryless, and connected minimizer manifolds). It is worth noting that the orientability of the manifold is not necessary in our setting. We define the volume measure $\mathcal{M}$ in a coordinate-free manner based on the Riemannian metric $g_{\mathcal{S}}$. Orientability, a stronger condition, ensures the existence of a global top-form rather than just a density, which is not required in our framework.
@@ Appendix B Proofs of Results in Section˜3
Consider the decreasing sublevel sets $E_{\varepsilon}:=\{x:\ g(x)-g_{\star}\leq\varepsilon\}$; if none is contained in $\mathcal{U}(\theta)$, compactness yields a limit point in $\mathcal{S}(\theta)\cap(\mathbb{R}^{d}\setminus\mathcal{U}(\theta))$, a contradiction.
@@
Fix $\theta\in\Theta$. By [73, Corollary 2.13] and the path-connectedness of $\mathcal{S}(\theta)$, the rank of $\partial_{x}^{2}g(\theta,x)$ is constant over $\mathcal{S}(\theta)$. Denote it by
@@
then the dimension of $\mathcal{S}(\theta)$ is $\texttt{k}(\theta)=d-\mathrm{r}(\theta)$.
@@
Moreover, implies a local Morse–Bott structure for $g(\theta,\cdot)$ [73, Corollary 2.13], hence the nonzero eigenvalues of $\partial_{x}^{2}g(\bar{\theta},x)$ are bounded away from 0 uniformly for $x$ near $\mathcal{S}(\bar{\theta})$. Therefore, for $n$ large enough, the perturbation above cannot change the number of nonzero eigenvalues, and thus
@@
Lemma B.3 ([9], Section 9.3).
@@
Proposition B.4 ([71], Lemma 6.4.7).
@@
For each fixed $\theta$, $\mathcal{S}(\theta)$ is a compact k-dimensional Riemannian manifold, hence $0<\mathcal{M}(\mathcal{S}(\theta))<\infty$ by the standard coordinate formula for the Riemannian volume element; see [53, Section “Riemannian volume”].
@@
To establish the boundedness of the density $\rho_{\theta}$, it suffices to show that $|\partial_{r}^{2}g(\theta,y(u,r))|$ is bounded. Since the condition ensures that $g(\theta,\cdot)$ is locally Morse-Bott (see [73, Section 2.3]), we have $\partial_{r}^{2}g(\theta,y(u,r))\big|_{r=0}\succeq\mu\cdot I_{(d-\texttt{k})\times(d-\texttt{k})}$.
@@ Appendix C Proofs of Results in Section˜4
$\mathrm{d}\mathcal{M}(x)\;=\;\sqrt{\det\bigl(\mathsf{g}_{ij}(x)\bigr)}\,\mathrm{d}x^{1}\cdots\mathrm{d}x^{\texttt{k}},$
@@
$\int_{U}f(x)\,\mathrm{d}\mathcal{M}(x)\;=\;\int_{U}f(x)\,\sqrt{\det\bigl(\mathsf{g}_{ij}(x)\bigr)}\,\mathrm{d}x^{1}\cdots\mathrm{d}x^{\texttt{k}},$
@@
Fix a point $p\in\mathcal{S}$ and $0<r<r_{0}$. By [71, Theorem 5.4.2], the exponential map
@@
$\mathcal{M}\bigl(\mathbb{B}_{\mathsf{g}}(p;r)\bigr)\;=\;\int_{\mathbb{B}_{\mathsf{g}}(p;r)}\mathrm{d}\mathcal{M}(x)\;=\;\int_{\mathbb{B}_{\texttt{k}}(0;r)}\bigl(\exp_{p}^{*}\mathrm{d}\mathcal{M}\bigr)(v),$ (39)
where $v=(v^{1},\dots,v^{\texttt{k}})$ are coordinates on $T_{p}\mathcal{S}\cong\mathbb{R}^{\texttt{k}}$ and $\mathrm{d}v$ is the standard Lebesgue measure on $\mathbb{R}^{\texttt{k}}$.
In normal coordinates at $p$, the exponential map $\exp_{p}$ is precisely the coordinate map, so $\exp_{p}^{*}\mathrm{d}\mathcal{M}$ has the expression
$(\exp_{p}^{*}\mathrm{d}\mathcal{M})(v)\;=\;\sqrt{\det\bigl(\mathsf{g}_{ij}(v)\bigr)}\,\mathrm{d}v^{1}\cdots\mathrm{d}v^{\texttt{k}},$
@@
$\mathcal{M}\bigl(\mathbb{B}_{\mathsf{g}}(p;r)\bigr)\;=\;\int_{\mathbb{B}_{\texttt{k}}(0;r)}\sqrt{\det\!\bigl(\mathsf{g}_{ij}(v)\bigr)}\;\mathrm{d}v.$ (40)
Step 2: Quadratic control of the metric in normal coordinates, uniformly in $p$. In a normal coordinate neighborhood of $p$, the metric admits the expansion (see, e.g., [71, Section 5.4])
@@
$\frac{3}{4}\,I\preceq\mathsf{g}(v)\preceq\frac{5}{4}\,I\,\Longrightarrow\,\Bigl(\frac{3}{4}\Bigr)^{\texttt{k}/2}\leq\sqrt{\det\bigl(\mathsf{g}_{ij}(v)\bigr)}\leq\Bigl(\frac{5}{4}\Bigr)^{\texttt{k}/2},\quad\text{for all }\|v\|\leq r_{*}.$ (44)
@@
$\Bigl(\frac{3}{4}\Bigr)^{\texttt{k}/2}{\texttt{Leb}}_{\texttt{k}}\bigl(\mathbb{B}_{\texttt{k}}(p;r)\bigr)\leq\mathcal{M}\bigl(\mathbb{B}_{\mathsf{g}}(p;r)\bigr)$
$=\int_{\mathbb{B}_{\texttt{k}}(0;r)}\sqrt{\det\bigl(\mathsf{g}_{ij}(v)\bigr)}\,\mathrm{d}v\leq\Bigl(\frac{5}{4}\Bigr)^{\texttt{k}/2}{\texttt{Leb}}_{\texttt{k}}\bigl(\mathbb{B}_{\texttt{k}}(p;r)\bigr).$
@@
where (a) follows from (45). Then, taking the minimum over $\beta$ at the r.h.s. of (46), we get $\min_{\beta^{\prime}}\phi_{0,\delta}(\theta,\beta^{\prime})-\min_{\beta}\phi_{\lambda,\delta}(\theta,\beta)\leq\frac{L_{f,1}}{\delta}(\mu^{\lambda}_{g}(\theta),\mu^{0}_{g}(\theta)))$. Interchanging $\min_{\beta}\phi_{\lambda,\delta}(\theta,\beta)$ and $\min_{\beta^{\prime}}\phi_{0,\delta}(\theta,\beta^{\prime})$, one obtains
@@ Appendix D Proofs of Results in Section˜5
$\int_{x:\|x\|>R}\|x\|^{n}e^{-c\|x\|^{2}}\mathrm{d}x=$
$\int_{0}^{\pi}\cdots\int_{0}^{\pi}\int_{0}^{2\pi}\int_{r>R}e^{-cr^{2}}r^{d+n-1}\sin^{d-2}(\phi_{1})\sin^{d-3}(\phi_{2})\ldots\sin(\phi_{d-2})\mathrm{d}r\mathrm{d}\phi_{1}\cdots\mathrm{d}\phi_{d-1}$
$\leq\int_{0}^{\pi}\cdots\int_{0}^{\pi}\int_{0}^{2\pi}\int_{r>R}e^{-cr^{2}}r^{d+n-1}\mathrm{d}r\mathrm{d}\phi_{1}\cdots\mathrm{d}\phi_{d-1}$
$\leq 2\pi^{d-1}\int_{R}^{+\infty}e^{-cr^{2}}r^{d+n-1}\mathrm{d}r\leq 2\pi^{d+n-1}c^{-\frac{d+n}{2}}\cdot\Gamma\left(\frac{d+n}{2},cR^{2}\right),$ (48)
where in the last inequality we use the definition of the upper incomplete Gamma function, i.e., $\Gamma(s,x):=\int_{x}^{+\infty}t^{s-1}e^{-t}\mathrm{d}t$. By Gautschi’s inequality [^26] By Gautschi-type’s inequality, $\Gamma(x,s)\leq sx^{s-1}e^{-x}$ for $x>s$ and $s\geq 1$. Note that for $s=1$, $\Gamma(s,x)=e^{-x}$., for $cR^{2}>(d+n)/2$ and $d+n\geq 2$,
$\int_{x:\|x\|>R}\|x\|^{n}e^{-c\|x\|^{2}}\mathrm{d}x\leq 2\pi^{d+n-1}\cdot c^{-\frac{d+n}{2}}\max\{(d+n)/2,1\}(cR^{2})^{\frac{d+n}{2}-1}e^{-cR^{2}}$
@@
$\mathbb{E}_{X\sim\mu_{\theta}^{\lambda}}[\|X\|^{n}\cdot\bm{1}\left[X\notin\mathbb{B}_{d}(0;2\mathsf{D})\right]]=\frac{1}{Z_{\lambda}}\int_{\|x\|\geq 2\mathsf{D}}\|x\|^{n}e^{-g(x)/\lambda}\mathrm{d}x,$
@@
$\mathbb{E}_{X\sim\mu_{\theta}^{\lambda}}[\|X\|^{n}\cdot\bm{1}\left[X\notin\mathbb{B}_{d}(0;2\mathsf{D})\right]]\leq\frac{1}{Z_{\lambda}}\int_{\|x\|\geq 2\mathsf{D}}\|x\|^{n}e^{-\frac{\mu_{\mathsf{qg}}(\|x\|-\mathsf{D})^{2}}{2\lambda}}\mathrm{d}x.$
@@
$\mathbb{E}_{X\sim\mu_{\theta}^{\lambda}}[\|X\|^{n}\cdot\bm{1}\left[X\notin\mathbb{B}_{d}(0;2\mathsf{D})\right]]\leq$
$\frac{1}{Z_{\lambda}}\cdot 2\pi^{d+n-1}\cdot\left(\frac{2\lambda}{\mu_{\mathsf{qg}}}\right)\cdot\max\{(d+n)/2,1\}\cdot\mathsf{D}^{d+n-2}e^{-\frac{\mathsf{D}^{2}}{\lambda}}.$ (51)
For a lower bound on $Z_{\lambda}$, we use the estimate (see [38, Remark A.21]): for a compact set $U\subset\mathbb{R}^{d}$ and $x_{0}\in\mathcal{S}$,
$\int_{U}e^{-g(x)/\lambda}\mathrm{d}x\geq\lambda^{\frac{d}{2}}\left[(2\pi)^{d/2}\det\left(\frac{\partial^{2}g(x_{0})}{\partial x^{2}}\right)^{-\frac{1}{2}}-c\lambda\right].$
@@
$\frac{1}{2}<\mathbb{P}_{X\sim\mu_{\theta}^{\lambda}}[f(\theta,X)<\beta^{*}(\theta)]\leq$
$\mathbb{P}[f(\theta,X)<\beta^{*}(\theta),X\in\mathbb{B}_{d}(0;2\mathsf{D})]+\mathbb{P}[f(\theta,X)<\beta^{*}(\theta),X\notin\mathbb{B}_{d}(0;2\mathsf{D})]$
@@
Given that (i) $\phi_{\lambda,\delta}(\theta,\cdot)$ is convex and Lipschitz, (ii) the feasible set is bounded, and (iii) the stochastic subgradient $\tilde{\partial}_{\beta}\phi_{\lambda,\delta}(\theta,\beta;X)$ satisfies $|\tilde{\partial}_{\beta}\phi_{\lambda,\delta}(\theta,\beta;X)|\leq\sigma:=1+\delta^{-1}$, we invoke the convergence guarantees for convex Lipschitz objectives [67, 51]. Accordingly, the output $\hat{\beta}$ of PSGD (see Algorithm˜1) satisfies
@@
after at most $\mathcal{O}(\mathrm{B}^{2}\sigma^{2}\epsilon_{v}^{-2}\log^{2}(\Delta^{-1}))$ iterations.
@@
$L\;\geq\;C\cdot\frac{\mathrm{B}^{2}}{\delta^{2}\epsilon_{v}^{2}}\cdot\log^{2}\left(\frac{1}{\epsilon_{v}\delta}\right).$
@@
By picking $\delta=\epsilon_{v}^{\texttt{k}}$ and $|I|\geq 4\tilde{C}_{f}\epsilon_{v}^{-2\texttt{k}-2}$ and
$L\;\geq\;\mathcal{O}\left(\texttt{k}^{2}\cdot\frac{\mathrm{B}^{2}}{\epsilon_{v}^{2(\texttt{k}+1)}}\cdot\log^{2}\left(\frac{1}{\epsilon_{v}}\right)\right),$
@@
Step 1: Standard zeroth-order analysis for the smoothed problem. We define the smoothed objective $\psi_{\rho}(\theta):=\mathbb{E}_{u\sim\text{uniform}(\mathbb{S}^{m-1})}[F_{\max}(\theta+\rho u)]$, where $u$ is uniformly distributed on the unit sphere $\mathbb{S}^{m-1}$. As shown in [68, Theorem 1 and Lemma 2], for any $L_{\mathsf{f},1}$-Lipschitz function $\mathsf{f}:\mathbb{R}^{m}\to\mathbb{R}$, the smoothed version $\mathsf{f}_{\rho}(\theta):=\mathbb{E}_{u\sim\text{uniform}(\mathbb{S}^{m-1})}[\mathsf{f}(\theta+\rho u)]$ satisfies:
@@
    $\nabla\mathsf{f}_{\rho}(\theta)\in\partial_{\rho}\mathsf{f}(\theta)$, the $\rho$-subdifferential of $\mathsf{f}$, as shown in [57, Theorem 3.1].
In our scenario, by using Section˜3.1, $F_{\max}$ is Lipschitz continuous with the constant $L_{F_{\max},1}:=L_{f,1}(1+L_{g,2}/\mu)$. By applying the aforementioned findings on $F_{\max}$, it follows that $\psi_{\rho}$ is $(L_{\psi,2}=\sqrt{m}L_{F_{\max},1}\rho^{-1})$-smooth. Furthermore, the gradient $\nabla\psi_{\rho}(\theta)$ is contained in the $\rho$-subdifferential set of $F_{\max}$ at the point $\theta$. Therefore, if a point $\theta$ satisfies $\|\mathcal{G}_{\Theta}(\theta,\nabla\psi_{\rho};\eta)\|\leq\epsilon$, then $\theta$ is an $(\epsilon,\rho,\eta)$-generalized Goldstein stationary point of $F_{\max}$ (see Section˜5.2). In what follows, we analyze the convergence of Algorithm 2 to a critical point of $\psi_{\rho}$ in terms of the gradient mapping.